Revise your calculations based the new information provided below and then answer the questions that follow. A company lends(b) Calculate the interest revenue to be recorded at the end of each year the note is outstanding. Interest revenue December(c) Prepare the journal entries to accrue the interest receivable for each year the note is outstanding. (If no entry is requ(d) Prepare the journal entry to record receiving the cash at the notes maturity. (If no entry is required for a transaction
Answer #1
a
Date General Journal Debit Credit
June 01 Notes Receivable 336000
      Cash 336000
b
Interest Revenue
December 31,Year 1 13720 =336000*7%*7/12
December 31,Year 2 23520 =336000*7%
December 31,Year 3 23520 =336000*7%
June 1,Year 4 9800 =336000*7%*5/12
c
Date General Journal Debit Credit
December 31 Interest Receivable 13720
      Interest Revenue 13720
December 31 Interest Receivable 23520
      Interest Revenue 23520
December 31 Interest Receivable 23520
      Interest Revenue 23520
d
Date General Journal Debit Credit
June 01 Cash 406560
      Interest Receivable 60760 =13720+23520+23520
      Interest Revenue 9800
      Notes Receivable 336000